### v1

# June 2, 2025

June 2, 2025
Muon (and Lepton) Anomalous Magnetic Moments
and Limits on Their Radii
Dimitri Bourilkov***e-mail: dimi@ufl.edu
University of Florida, Gainesville,
Florida, 32611, USA
The limits on lepton radii and contact interaction scales are used to derive limits on the deviations of lepton anomalous magnetic dipole moments (g-2)/2 from the Standard Model predictions. In the case of quadratic dependence of the deviations on the lepton mass the limits for electrons are better compared to low energy measurements, for muons somewhat weaker than the latest precision experiments, and for taus by far the best.
Without a doubt the Standard Model (SM) is very successful in describing the precision measurements performed at the highest energy accelerators. The SM assumes that fundamental particles, like the leptons, are point-like.
At the energy frontier, the lepton-pair measurements at LEP at energies above the Z resonance mass provide very sensitive probes of the point-like structure of the electron [1, 2], the muon and the tau [3].
At the precision frontier, the Fermilab Muon g-2 experiment [4] has generated excitement when comparing their measurements to the best available theory predictions. If the anomalous magnetic moments deviate from theory, this may be a sign for lepton substructure and finite lepton radii, or for new interactions. In this study we analyze the interplay between lepton radii, contact interactions and anomalous magnetic moments.
The high precision measurements of the magnetic dipole moment $\rm(g-2)_{e}$ of the electron have been used to put stringent limits on the electron radius $\rm r_{e}$ [5, 6]. If non-standard contributions to $\rm(g-2)_{e}$ scale linearly with the electron mass, the bound is of the order of $\rm r_{e}\sim 10^{-23}\ m$. As a natural consequence of chiral symmetry, these corrections are expected to scale quadratically with the lepton mass [5]. In this case the bound is reduced to $\rm r_{e}\sim 10^{-18}\ m$.
The L3 collaboration has analyzed the fermion-pair measurements above the Z pole for deviations from the point-like structure [3]. A combined analysis of the electron data is also available [1, 2]. The SM cross sections for the reactions $\mathrm{e}^{+}\mathrm{e}^{-}\rightarrow f\bar{f}$ are modified as follows:
$\frac{d\sigma}{dq^{2}}=\left(\frac{d\sigma}{dq^{2}}\right)_{SM}F_{e}^{2}(q^{2})F_{f}^{2}(q^{2}),$ (1)
where $q^{2}$ is the Mandelstam variable $s$ or $t$ for $s$– or $t$–channel exchange, the latter only important for Bhabha scattering, and the form factors of the initial and final state fermions are denoted as $F_{e}$ and $F_{f}$, respectively. They are parameterized by a Dirac form factor:
$F(q^{2})=1+\frac{1}{6}q^{2}r^{2},$ (2)
where $r$ is the radius of the fermions.
The upper limits on the fermion radii obtained from the LEP data are shown in Table 1. They are derived with the assumption $F_{e}=F_{f}$. For the $\mathrm{\mu^{+}\mu^{-}}$ and $\mathrm{\tau^{+}\tau^{-}}$ final states the limits given in Table 1 will increase by a factor of $\sqrt{2}$ under the most conservative assumption that the electron is point-like ($F_{e}\equiv 1$).
  Channel    $r$ [m]    $\delta a_{l}$
$\mathrm{e^{+}e^{-}}$ $\rm 2.8\cdot 10^{-19}$ $\rm 5.3\cdot 10^{-13}$
$\mathrm{\mu^{+}\mu^{-}}$ $\rm 2.4\cdot 10^{-19}$ $\rm 1.7\cdot 10^{-8}$
$\mathrm{\tau^{+}\tau^{-}}$ $\rm 4.0\cdot 10^{-19}$ $\rm 1.3\cdot 10^{-5}$
Table 1: Upper limits on the fermion radii at 95 % confidence level from form factors for electrons [1], muons and taus [3], and corresponding values for $\delta a_{l}$.
From the combined analysis[1] of the LEP collaborations data on Bhabha scattering the best upper limit on the electron radius at 95% confidence level was derived:
$\rm r_{e}<2.8\cdot 10^{-19}\ m.$ (3)
This limit is substantially lower than the limit derived from $\rm(g-2)_{e}$ measurements in the case where the deviations from the SM of the magnetic dipole moment of the electron depend quadratically on its mass.
To connect the bounds on lepton radii to deviations for the anomalous magnetic moment we will use a quadratic dependence on the masses:
$\delta a_{l}=m_{l}^{2}r_{l}^{2}$ (4)
where $\delta a_{l}$ is the deviation of the anomalous magnetic dipole measurement from the SM prediction for lepton l, $m_{l}$ is the lepton mass in GeV, and $r_{l}$ is the lepton radius in units 1/GeV. To help the reader we convert most distances to meters.
In searches for new interactions beyond the SM, a standard framework is the most general combination of helicity conserving dimension-6 operators [7]. In this scheme, we have as parameters a coupling strength, $g$, and an energy scale, $\Lambda$, which can be viewed as the scale of compositeness. At energies well below $\Lambda$, this results in an effective Lagrangian leading to four-fermion contact interactions.
As the LEP experiments have observed no statistically significant deviations from the SM, they have derived one-sided lower limits on the scale $\rm\Lambda$ of contact interactions at 95% confidence level [8] under the assumption of a strong coupling $\rm\frac{g^{2}}{4\pi}=1$. The best limits are summarized in Table 2.
  Channel     ${}~{}~{}\Lambda_{VV}^{+}$    $r$    $\lambda_{VV}^{+}$    $r$
   [TeV]    [m]    [TeV]    [m]
$\mathrm{e^{+}e^{-}}$ 21.3 $\rm 9.2\cdot 10^{-21}$ 1.9 $\rm 1.0\cdot 10^{-19}$
$\mathrm{\mu^{+}\mu^{-}}$ 18.9 $\rm 1.0\cdot 10^{-20}$ 1.7 $\rm 1.2\cdot 10^{-19}$
$\mathrm{\tau^{+}\tau^{-}}$ 15.8 $\rm 1.2\cdot 10^{-20}$ 1.4 $\rm 1.4\cdot 10^{-19}$
Table 2: Lower limits at 95 % confidence level for electrons [2], muons and taus [8] on the contact interaction scales $\Lambda$ for the VV model with positive interference, and for the corresponding “QED rescaled” scales $\lambda$. The scales are converted to upper limits on the fermion radii.
How to compare the limits on lepton radii to the limits on the contact interaction scale? They are different parts of the same puzzle. For example, the limit on the electron size $\rm r_{e}<2.8\cdot 10^{-19}\ m$ translates to a mass scale $\rm M>0.7\ TeV$.
In contrast, for contact interactions very stringent limits $\sim$ 15-20 TeV were derived. By convention, a strong coupling $\rm\frac{g^{2}}{4\pi}=1$ for the novel interactions is postulated, which may be too optimistic. In a more conservative approach, if we assume an electromagnetic strength coupling, the limits are reduced substantially:
$\rm\lambda=\sqrt{\alpha_{QED}}\cdot\Lambda=0.0884\cdot\Lambda$ (5)
where $\alpha_{QED}$ at the Z resonance is used and small effects, at the percent level, of the running of $\rm\alpha_{QED}$ to higher energies are ignored. $\lambda$ is the “QED modified”, lower, scale. A careful study shows that the VV contact interaction model with positive interference modifies the differential cross section in a similar way as a form factor, so we will use these results for comparisons. The results for $\Lambda$ and $\lambda$ are compiled in Table 2. As can be seen, in the case of $\lambda$ the bounds on the lepton radii are close to the form factor bounds.
The limits on $\delta a_{l}$ coming from form factors are summarized in Table 1, and from contact interactions for $\lambda$ in Table 3. Under the quadratic dependence assumption, our electron limits are better than the (g-2) measurements.
The situation is reversed for muons, where the latest comparison [4, 9] gives
$\delta a_{\mu}=2.6\pm 6.6\cdot 10^{-10}$ (6)
compared to our limits
$\delta a_{\mu}=1.7\cdot 10^{-8}\ \ \rm{for\ form\ factor}$ (7)
$\delta a_{\mu}=4.1\cdot 10^{-9}\ \ \rm{for\ contact\ interaction}.$ (8)
If a novel interaction happens to exceed electromagnetic strength, the limit can become more competitive.
The situation is reversed for tau leptons: as experimental measurements of $\delta a_{\tau}$ are very difficult, our limits by far exceed direct approaches.
  Channel    $r$ [m]    $\delta a_{l}$
$\mathrm{e^{+}e^{-}}$ $\rm 1.0\cdot 10^{-19}$ $\rm 6.7\cdot 10^{-14}$
$\mathrm{\mu^{+}\mu^{-}}$ $\rm 1.2\cdot 10^{-19}$ $\rm 4.1\cdot 10^{-9}$
$\mathrm{\tau^{+}\tau^{-}}$ $\rm 1.4\cdot 10^{-19}$ $\rm 1.6\cdot 10^{-6}$
Table 3: Upper limits on the fermion radii at 95 % confidence level from contact interactions with scale $\lambda$ for electrons, muons and taus, and corresponding values for $\delta a_{l}$
To conclude: we used the limits on lepton radii and contact interaction scales to derive limits on the deviations of lepton anomalous magnetic dipole moments (g-2)/2 from the Standard Model predictions. In the case of quadratic dependence of the deviations on the lepton mass the limits for electrons are better compared to low energy measurements, for muons somewhat weaker than the latest precision experiments, and for taus by far the best.
Acknowledgments
The author is grateful to the University of Florida High Energy Group for supporting this type of work.